OpenStudy (anonymous):

The product of two consecutive integers is 420. An equation is written in standard form to solve for the smaller integer by factoring. What is the constant of the quadratic expression in this equation? x2 + x +.... = 0

OpenStudy (anonymous):

@Nnesha @robtobey

hartnn (hartnn):

2 consecutive integers are x,x+1 their product is 420

hartnn (hartnn):

x(x+1) =420 try an simplify this... bring 420 on left side

OpenStudy (anonymous):

so 420x(x-1) ?

hartnn (hartnn):

*sigh* whats x(x+1) equal to?? use distributive property...

OpenStudy (anonymous):

1

hartnn (hartnn):

wow ok, here are the steps, go through them and ask doubts :) \(x(x+1) =420 \\x^2 +x = 420 \\ x^2+x-420 = 420-420 \\ x^2+x-420 =0\) so your answer is -420

OpenStudy (anonymous):

@hartnn Im Not The Best At Math Sorry :(

hartnn (hartnn):

don't be sorry! practice and become good at it :D

OpenStudy (anonymous):

So The Answer is -420

hartnn (hartnn):

yes